# Zero-Shot Cross-Lingual NER Using Phonemic Representations for Low-Resource Languages

### v1

@@ Abstract
Existing zero-shot cross-lingual NER approaches require substantial prior knowledge of the target language, which is impractical for low-resource languages. In this paper, we propose a novel approach to NER using phonemic representation based on the International Phonetic Alphabet (IPA) to bridge the gap between representations of different languages. Our experiments show that our method significantly outperforms baseline models in extremely low-resource languages, with the highest average F-1 score (46.38%) and lowest standard deviation (12.67), particularly demonstrating its robustness with non-Latin scripts.
@@ 2.2 Phonemic Representation
Phonological traits of languages are useful in understanding different languages, as they often share similar pronunciations for similar entities. It is particularly beneficial for NER, where many items, such as geopolitical entities and personal names, are pronounced similarly across various languages. While phonological information has been shown to be helpful in language understanding for cross-lingual transfer Chaudhary et al. (2018); Sun et al. (2021); Bharadwaj et al. (2016), it has not been explored as a standalone representation for NER, especially on low-resource languages. Given that creating rule-based transcription module for most low-resource languages takes only a few hours and limited training, we use IPA to enable zero-shot cross-lingual NER on languages with very scarce resources, without requiring any additional corpus for those languages.
@@ 4.2 Baseline Models
We use mBERT Devlin et al. (2019) and CANINE Clark et al. (2022), both grapheme-based language models, as baselines to compare to XPhoneBERT Nguyen et al. (2023), a phoneme-based language model. All three models are BERT-like transformer architectures pre-trained on a Wikipedia corpora of multiple languages: mBERT and CANINE are trained on the same 104 languages, while XPhoneBERT is trained on 94 languages and locales. Initializing with pre-trained weights from Huggingface [^2] https://huggingface.co/, we train the encoders with a fully connected layer added at the end of each encoder for NER prediction.
@@ 5.2 Zero-Shot NER on Unseen Languages
Figure 3 shows a qualitative result of zero-shot inference on Sinhala, a language that is not in the pre-training data any model. While the character-based model (a) fails to generalize to the language with different writing system, the phoneme-based model (b) successfully predicts the named entity tags due to the similar pronunciation of “China” and “Russia” across the languages. These results indicate the robustness provided by phonemic representations, validating our hypothesis about the advantages they convey in NER tasks.
@@ 5.3 Robustness Across Writing Systems
One of the important advantages of using phonemic representations for named entity recognition is that it allows use of IPA. Using IPA for multilingual tasks provides a unified notation system. Observing the significant performance drop of mBERT on unseen low-resource languages (Figure 2), we consider this gap is largely attributed to the different writing systems of languages. Figure 4 shows the distribution of F1 scores of each model on Latin and non-Latin languages from Case 1. mBERT, which performs the strongest on seen languages, exhibits the largest performance discrepancy between Latin and non-Latin based languages when evaluated on unseen languages. This highlights the limitation of the grapheme-based model, as it depends on the specific scripts.
On the other hand, the phoneme-based model—XPhoneBERT—demonstrates the most consistent performance over different unseen languages with little performance gap between Latin-based and non-Latin-based languages. This suggests that taking advantage of phonemes with its unified notation system allows for better generalization on extremely low-resource languages.
Figure 4: Performance distribution of each model on Latin and non-Latin languages from unseen languages.
@@ 8 Ethics Statement
Appendix A Appendix
A.1 Implementation Details
@@
A.2 Quantitative Results of Case 2 and Case 3
@@
grapheme mBERT 10.71 44.76 38.48 55.07 18.7 12.58 62.37 79.51 40.27 25
grapheme CANINE 26.31 43.35 51.3 59.48 27.19 22.38 54.74 80.7 45.68 19.99
phoneme (ours) XPhoneBERT 43.61 38.91 38.07 51.9 44.82 31.03 49.67 73.05 46.38 12.67
@@
A.3 Comparison of Latin and Non-Latin Languages
In Figure 5, we visualize the results of the experiment separately for Latin and non-Latin languages in all cases. Compared to mBERT and CANINE that exhibit significant performance gaps between Latin and non-Latin languages, XPhoneBERT shows little difference in performance distribution.
Figure 5: Latin and non-latin comparison
A.4 Language codes
In Table 4, we organized both ISO 639-1 and ISO 639-3 languages codes of all the languages used in the experiments.
Lang Code
ISO 639-1 ISO 639-3
Amharic am amh
Assyrian Neo-Aramaic aii aii
Ayacucho quechua qu quy
Cebuano ceb ceb
Croatian hr hrv
English en eng
Esperanto eo epo
Ilocano ilo ilo
Javanese jv jav
Khmer km khm
Kinyarwanda rw kin
Korean ko kor
Kyrgyz ky kir
Malay ms msa
Malayalam ml mal
Maltese mt mlt
Maori mi mri
Marathi mr mar
Punjabi pa pan
Sinhala si sin
Somali so som
Spanish es spa
Tajik tg tgk
Telugu te tel
Turkmen tk tuk
Urdu ur urd
Uyghur ug uig
Uzbek uz uzb
Yoruba yo yor
Table 4: Language codes for all the languages used in the experiments.
A.5 Benchmark and License.
@@ A.4 Language codes
Dataset Lang. Train Dev Test License
WikiANN eng 20k 10k 10k ODC-BY
sin 100 100 100
som 100 100 100
mri 100 100 100
quy 100 100 100
uig 100 100 100
aii 100 100 100
kin 100 100 100
ilo 100 100 100
epo 15k 10k 10k
khm 100 100 100
tuk 100 100 100
amh 100 100 100
mlt 100 100 100
ori 100 100 100
san 100 100 100
ina 100 100 100
grn 100 100 100
bel 15k 1k 1k
kur 100 100 100
snd 100 100 100
tgk 100 100 100
yor 100 100 100
mar 5k 1k 1k
jav 100 100 100
urd 20k 1k 1k
msa 20k 1k 1k
ceb 100 100 100
hrv 20k 10k 10k
mal 10k 1k 1k
tel 1k 1k 1k
uzb 1k 1k 1k
pan 100 100 100
kir 100 100 100
Table 5: Statistics and license types for the dataset. The table lists the number of examples in the training, development, and testing sets for languages in the WikiANN dataset. The dataset is strictly used within the bounds of these licenses.

### v2

@@ Abstract
Existing zero-shot cross-lingual NER approaches require substantial prior knowledge of the target language, which is impractical for low-resource languages. In this paper, we propose a novel approach to NER using phonemic representation based on the International Phonetic Alphabet (IPA) to bridge the gap between representations of different languages. Our experiments show that our method significantly outperforms baseline models in extremely low-resource languages, with the highest average F1 score (46.38%) and lowest standard deviation (12.67), particularly demonstrating its robustness with non-Latin scripts. Our codes are available at https://github.com/Gabriel819/zeroshot_ner.git
@@ 2.2 Phonemic Representation
Phonological traits of languages are useful in understanding different languages, as they often share similar pronunciations for similar entities. It is particularly beneficial for NER, where many items, such as geopolitical entities and personal names, are pronounced similarly across various languages. Moreover, phonemes are represented in IPA, which is shared across all languages. By providing a universal script, phonemic representations help the model better address low-resource languages, as the poor NER performance in these languages comes significantly from the model’s limited ability to handle relevant scripts Muller et al. (2020); Severini et al. (2022).
While phonological information has been shown to be helpful in language understanding for cross-lingual transfer Chaudhary et al. (2018); Sun et al. (2021); Bharadwaj et al. (2016); Leong and Whitenack (2022), few works have explored its benefits compared to orthographic input, particularly in zero-shot scenario where the target language is not available for fine-tuning. Given that creating rule-based transcription module for most low-resource languages takes only a few hours and limited training, we use IPA to enable zero-shot cross-lingual NER on languages with very scarce resources, without requiring any additional corpus for those languages.
@@ 4 Experiments
Here we provide information about the datasets and models we employed for the experiments. More implementation details including hyperparameters are provided in Appendix A.
@@ 4.2 Baseline Models
We use mBERT Devlin et al. (2019) and CANINE Clark et al. (2022), both grapheme-based language models, as baselines to compare to XPhoneBERT Nguyen et al. (2023), a phoneme-based language model. All three models are BERT-like transformer architectures pre-trained on a Wikipedia corpora of multiple languages: mBERT and CANINE are trained on the same 104 languages, while XPhoneBERT is trained on 94 languages and locales. Initializing with pre-trained weights from Huggingface [^2] https://huggingface.co/, we train the encoders with a fully connected layer added at the end of each encoder for NER prediction. We also provide a comparison with XLM-R Conneau et al. (2020) in Appendix D, with different set of languages for Case 1, Case 2, and Case 3.
@@ 5.2 Zero-Shot NER on Unseen Languages
Figure 3 shows a qualitative result of zero-shot inference on Sinhala, a language that is not in the pre-training data of any model. While the character-based model (a) fails to generalize to the language with different writing system, the phoneme-based model (b) successfully predicts the named entity tags due to the similar pronunciation of “China” and “Russia” across the languages. These results indicate the robustness provided by phonemic representations, validating our hypothesis about the advantages they convey in NER tasks.
Figure 4: Performance distribution of each model on languages using Latin and non-Latin scripts from unseen languages.
@@ 5.3 Robustness Across Writing Systems
One of the important advantages of using phonemic representations for named entity recognition is that it allows use of IPA. Using IPA for multilingual tasks provides a unified notation system. Observing the significant performance drop of mBERT on unseen low-resource languages (Figure 2), we consider this gap is largely attributed to the different writing systems of languages. Figure 4 shows the distribution of F1 scores of each model on languages using Latin and non-Latin scripts from Case 1. mBERT, which performs the strongest on seen languages, exhibits the largest performance discrepancy between languages that use Latin and non-Latin scripts when evaluated on unseen languages. This highlights the limitation of the grapheme-based model, as it depends on the specific scripts.
On the other hand, the phoneme-based model—XPhoneBERT—demonstrates the most consistent performance over different unseen languages with little performance gap between Latin and non-Latin scripts. This suggests that taking advantage of phonemes with its unified notation system allows for better generalization on extremely low-resource languages.
@@ 8 Ethics Statement
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2022-00143911,AI Excellence Global Innovative Leader Education Program)(90%) and ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT)(RS-2020-II201819)(10%).
Appendix A Implementation Details
@@
Appendix B Quantitative Results of Case 2 and Case 3
@@
Figure 5: Performance distribution of each model on languages using Latin and non-Latin scripts.
Appendix C Comparison of Latin and Non-Latin Languages
Figure 5 shows the performance distribution of zero-shot NER on all languages, where the performance distribution of languages using Latin and non-Latin scripts are visualized separately. Compared to mBERT and CANINE that exhibit significant performance gaps between Latin and non-Latin scripts, XPhoneBERT shows little difference in performance distribution.
Case Input Model Languages AVG STD
CASE 1 mri quy aii kin ilo tgk yor ceb
grapheme XLM-R 29.93 65.09 14.58 49.05 71.29 40.96 55.01 64.31 48.78 19.42
phoneme (ours) XPhoneBERT 38.07 51.90 31.03 49.67 73.05 48.93 50.87 58.72 50.28 12.62
CASE 2 tuk mlt ina grn
grapheme XLM-R 53.63 58.81 72.55 44.80 57.45 11.61
phoneme (ours) XPhoneBERT 61.17 52.58 68.68 49.95 58.10 8.53
CASE 3 epo khm amh ori san bel kur snd sin som uig
grapheme XLM-R 73.82 48.85 64.89 54.93 59.92 82.93 75.61 42.90 64.91 55.52 61.74 62.37 11.88
phoneme (ours) XPhoneBERT 75.26 31.86 44.85 40.73 59.42 77.61 52.95 47.28 43.61 38.91 44.82 50.66 14.60
Table 4: Zero-shot F1 score (%) result in Case 1, 2, and 3 languages specifically selected for XLM-R and XPhoneBERT.
Appendix D Comparison with XLM-R
Here we provide additional results of zero-shot inference of XLM-R(base) Conneau et al. (2020) and XPhoneBERT Nguyen et al. (2023). XLM-R shares the training objective with XPhoneBERT, removing the variance that comes from the difference in pre-training objectives. Since XLM-R is trained on different set of languages, we listed up another sets of languages that corresponds to Case1, Case2, and Case3, which refers to the set of languages not trained on both models, languages trained on XPhoneBERT not on XLM-R, and languages trained on XLM-R not on XPhoneBERT, respectively.
As shown in Table 4, the result for Case1 is consistent with our findings in the main experiments, exhibiting relatively stable performance with XPhoneBERT (higher average and lower standard deviation). Also, languages that are written in non-Latin scripts perform significantly better with XPhoneBERT, which aligns with our analysis in the main text as well.
Appendix E Language Codes
In Table 5, we have listed ISO 639-3 language codes of all languages used in the experiments.
Language ISO 639-3
Amharic amh
Assyrian Neo-Aramaic aii
Ayacucho quechua quy
Cebuano ceb
Croatian hrv
English eng
Esperanto epo
Ilocano ilo
Javanese jav
Khmer khm
Kinyarwanda kin
Korean kor
Kyrgyz kir
Malay msa
Malayalam mal
Maltese mlt
Maori mri
Marathi mar
Punjabi pan
Sinhala sin
Somali som
Spanish spa
Tajik tgk
Telugu tel
Turkmen tuk
Urdu urd
Uyghur uig
Uzbek uzb
Yoruba yor
Table 5: Language codes for all languages used in the experiments.
Appendix F Benchmark and License
@@
Dataset Lang. Script Train Dev Test License
WikiANN eng Latn 20k 10k 10k ODC-BY
sin Sinh 100 100 100
som Latn 100 100 100
mri Latn 100 100 100
quy Latn 100 100 100
uig Arab 100 100 100
aii Syrc 100 100 100
kin Latn 100 100 100
ilo Latn 100 100 100
epo Latn 15k 10k 10k
khm Khmr 100 100 100
tuk Latn 100 100 100
amh Ethi 100 100 100
mlt Latn 100 100 100
ori Orya 100 100 100
san Deva 100 100 100
ina Latn 100 100 100
grn Latn 100 100 100
bel Cyrl 15k 1k 1k
kur Latn 100 100 100
snd Arab 100 100 100
tgk Cyrl 100 100 100
yor Latn 100 100 100
mar Deva 5k 1k 1k
jav Latn 100 100 100
urd Arab 20k 1k 1k
msa Latn 20k 1k 1k
ceb Latn 100 100 100
hrv Latn 20k 10k 10k
mal Mlym 10k 1k 1k
tel Telu 1k 1k 1k
uzb Cyrl 1k 1k 1k
pan Guru 100 100 100
kir Latn 100 100 100
Table 6: Statistics and license types for the dataset. The table lists the script, number of examples in the training, development, and testing sets for languages in the WikiANN dataset. The dataset is strictly used within the bounds of these licenses.